OpenStudy (anonymous):

The Surface areas of two similar soilds are 225m^2 and 1089m^2. The volume of the larger one is 1331m^3. What is the volume of the smaller one?

OpenStudy (anonymous):

Ratio surface/surface is the square of side ratio Ratio S/s = 1089/225 = 4.84 = (2.2)^2 Ratio volume V/v is the cube of side ratio. Ratio V/v = (2.2)^3 = 10.648 Volume of the small solid: v = V/10.648 = 1331/10.648 = 125 m^3